OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 4x – y = 4 x + 4y = 18

OpenStudy (anonymous):

(1, 4) / (4, 2) / (4, 1) / (2, 4)

OpenStudy (anonymous):

what have you come up with so far?

OpenStudy (anonymous):

i got (4, 1) ?

OpenStudy (anonymous):

just solve for whatever you think would be easier. You can solve for y in the first equation or x in the second equation. Personally i think it'd be easier to solve for x in the second equation. We can solve for x in the second equation by subtracting 4y from both sides x = 18 - 4y Now that we know that x equals, we can plug it into the first equation. We will plug in ALL of these numbers into x in the first equation which is 4x - y = 4 so, plug in the entire 18-4y into x. 4 (18-4y) - y = 4 we have to use the distributive property. 4 times 18 equals 72, 4 times negative 4y equals -16y. 72 - 16y - y = 4 let's combine like terms 72 - 17y = 4 subtract 72 from both sides -17y = -68 divide by 17 to solve for y y = 4 NOW that we know what y equals, let's plug it back into the equation where we found out what x equals, which is x = 18 - 4y plug in 4 x = 18 - 4(4) x = 18 - 16 x = 2 your solution is (2,4)

OpenStudy (anonymous):

:( i put thought into that :"(

OpenStudy (anonymous):

thank you! lol, your really helpful!

OpenStudy (anonymous):

thanks :")